Find three consecutive integers if twice the middle integer is equal to the sum of the first and the third.
Answer by checkley71(8403)

x,(x+1),(x+2) are the 3 integers.
2(x+1)=x+(x+2)
2x+2=2x+2
0=0 thus the first interger is=0
proof
2(0+1)=0+0+2
20+2=2
2=2
--------------------------
1,2 & 3 also satisfies this problem
2*2=1+3
4=4